ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

2. Algebra

1991 MS Classification: 20E07, 20E22, 20E10 Vahagn, Mikaelian, Yerevan State University, Armenia Embeddings of countable groups into 2-generator groups with additional properties The well known theorem of Higman, B. Neumann and P. Neumann on embeddability of an arbirary countable group H into a 2-generator group G can be strongly generalised. For an arbitrary non-trivial word set V ⊆ F_{∞} the 2-generator group G can be chosen su that the image of H under this embedding is subnormal in G and lies in the verbal subgroup V (G). The method developed gives generalizations of some related results on embeddings of solvable groups into 2-generator solvable groups, etc. .. References V. H. Mikaelian, Subnormal embedding theorems for groups, (to appear) H. Heineken, V. H. Mikaelian, Normal verbal embeddings of groups, (to appear).